Data on the compressive strength of a concrete block after testing are as follows: 2.54 MPa, 2.59 MPa, 2.50 MPa, 2.56 MPa, and 2.52


MPa. Find:


e. 1


the average compressive strength


1
Expert's answer
2020-10-05T10:54:13-0400

"\sigma_{avr}=\frac{2.54+2.59+2.50+2.56+2.52}{5}=2.54(MPa)"